# supplement.pdf

### v1

@@
Abstract
In this paper, we present TexPro, a novel method for high-fidelity material generation for input 3D meshes given text prompts. Unlike existing text-conditioned texture generation methods that typically generate RGB textures with baked lighting, TexPro is able to produce diverse texture maps via procedural material modeling, which enables physical-based rendering, relighting, and additional benefits inherent to procedural materials. Specifically, we first generate multi-view reference images given the input textual prompt by employing the latest text-to-image model. We then derive texture maps through a rendering-based optimization with recent differentiable procedural materials. To this end, we design several techniques to handle the misalignment between the generated multi-view images and 3D meshes, and introduce a novel material agent that enhances material classification and matching by exploring both part-level understanding and object-aware material reasoning. Experiments demonstrate the superiority of the proposed method over existing SOTAs and its capability of relighting.
Keywords: 3D asset creation, text-guided texturing, procedural material, MLLM
footnotetext: *Authors contributed equally.footnotetext: †Corresponding author.
@@ 1 Introduction
3D mesh texturing is a critical process in the realm of 3D modeling and visualization, which holds significant importance across various applications such as AR/VR, gaming, filmmaking, etc. By transforming a basic geometric shape into a visually detailed and realistic object, mesh texturing not only improves the aesthetic appeal of the model but also greatly enhances immersion and realism in digital environments. To achieve photorealistic rendering that responds accurately to various lighting conditions, we need to generate object material that includes a set of texture maps like diffuse and specular reflections, glossiness, surface roughness, etc. As a result, traditional mesh texturing normally involves substantial manual effort. Recent advances in 2D image generation have spurred the development of automated mesh texturing methods [6, 33], but they normally generate RGB textures with baked lighting, which limits their application in downstream tasks.
@@
In this paper, we propose a novel method for producing high-quality materials for 3D meshes based on user-provided text prompts as shown in Fig. 1. Unlike previous methods that directly generate pre-baked 2D texture maps, our approach is able to produce diverse texture maps via procedural material modeling, which enables physical-based rendering, relighting, and additional benefits inherent to procedural materials, such as resolution independence (i.e., producing textures at any desired resolution), re-editability, efficient storage, and so on.
@@
Specifically, we first generate multi-view reference images given the input textual prompt by employing the latest text-to-image models [34, 43], and then obtain diverse texture maps through a rendering-based optimization with recent differentiable procedural materials [35, 24]. After the optimization, we can essentially obtain procedural materials for the mesh, thus supporting downstream applications. However, it is non-trivial to design such a system.
@@
First, despite using additional rendered depth or normal maps as conditions [43], the generated images from the text-to-image model (e.g., Stable Diffusion [34]) often fail to align accurately with the meshes and may not maintain consistency across different views. To handle these misalignment challenges, as shown in Fig. 2, we propose to exploit an improved segmentation method Matcher [27] to extract the masks of corresponding objects in generated images, which will be further smoothed and integrated with the rendered masks to get the final aligned masks. Additionally, we design an adaptive camera sampling method that ensures each material part exceeds a specified pixel count threshold and assigns a unique material to each part to address the inconsistency across views.
@@
Second, an initial estimation of material properties is requisite for differentiable procedural material rendering. However, the presence of baked lighting in generated images and the confusion of different materials with similar color distributions make initial material classification and matching challenging for classical methods [31, 15]. To overcome these challenges, we leverage the capabilities of Multi-modal Large Language Models (MLLMs) to create a novel material agent. Given the masked prompt image, the material agent is able to provide the possible material types in the order of likelihood for the corresponding object via the part-level understanding and object-aware material reasoning, which further facilitates the rendering-based optimization with procedural materials.
@@
Additionally, to better recover the material properties, we also consider the environmental lighting during the optimization. To make the optimization tractable, we adopt a vanilla lighting setup to fit the lighting environment in the image. Given the object mesh, our method adaptively creates the floor and four walls in the scene of the differentiable renderer, and adaptively places initial area lights on each wall and above the object. Then we optimize the RGB intensity of these lights together with procedural materials.
@@
  • We propose a novel framework that is able to create high-fidelity textures for input meshes given text prompts and enable physical-based rendering (PBR) and relighting.
  • In order to handle the misalignment between the generated multi-view images and 3D meshes, we propose to combine the rendered mask with an advanced segmentation method, i.e., Matcher [27], to achieve accurately aligned part-level masks, and design an adaptive camera sampling strategy to deal with inconsistency across views.
@@
  • The experiments show that the proposed method outperforms existing SOTAs and maintains the capabilities of relighting.
@@ 2 Related Work
Texture generation. Texture maps are essential for 3D geometric models, which involve mapping from pre-defined 2D planar images onto 3D models, endowing 3D models with vital color information. The traditional texture creation involves cumbersome manual drawing, assembling repetitive patterns, or stitching multi-view images [38, 3]. With the development of increasingly high-quality 3D datasets and advancements in text-to-image generative techniques  [32, 34, 43], learning-based approaches have been proposed to generate high-quality textures [20, 28, 6, 33, 40]. Text2Tex [6] incorporates a depth-aware image inpainting model to synthesize high-resolution partial textures progressively from multiple viewpoints, which has also been widely employed in text-to-scene tasks [9, 14, 42]. Similarly, TEXTure [33] also applies an iterative scheme to paint 3D models, while presenting a trimap representation and novel elaborated diffusion sampling process to tackle inconsistencies issues. Although existing methods are capable of generating high-quality textures, they are limited to generating only diffuse maps with baked lighting, lacking support for relighting. Additionally, the generated textures often suffer from Janus (multi-faced) problem [2] and lack realism.
Procedural material modeling and capture. Procedural material has been the standard of material modeling in industry. The materials are represented as node graphs denoted with simple image processing operations which are combined to produce real-world spatially varying BRDF [5] material maps. Given user-specified images or text prompts, many approaches [16, 35, 17, 12, 24] have been proposed to recover the parameters of a predefined procedural material graph to align with the input. Moreover, Hu et al. [18] moved beyond parameter regression, utilizing the diffusion model to generate graph structures from text or image conditions. It is worth noting that, MATch [35] proposes a differentiable procedural material method that translates the procedural graphs into differentiable node graphs, enabling single-shot high-quality procedural material capture.
Differentiable rendering. Differentiable rendering has consistently remained a significant research topic in computer graphics and vision. The gradient in rendering is required with respect to numerous factors, and specially several differentiable renderers widely used in community like Redner [25], Mitsuba 2/3 [30, 19], Nvdiffrast [23] and PSDR-CUDA [41] have been developed to address challenges arising from the non-differentiable terms in rendering integral. We leverage the differentiable rendering to backpropagate the gradient from pixel space to the parameters of procedural materials, thereby optimizing the materials via gradient descent.
@@ 3 Method
Most 3D meshes from online resource repositories like BlenderKit or Sketchfab, as well as meshes from existing 3D datasets [10, 26, 37, 8, 7], are segmented into different material parts by their authors during creation, e.g., a chair consists of the frame, backrest and cushion. Moreover, part-level segmentation on a complete mesh can be achieved through shape analysis methods [13, 22, 39]. As a result, given textual prompts, we aim to generate appropriate materials for each material part of the input 3D mesh model and achieve photorealistic and relightable texturing.
@@
As shown in Fig. 2, our method textures the mesh with PBR materials through three steps: multi-view reference images generation, materials initialization under material agent guidance, and materials optimization using differentiable rendering. In this section, we begin by introducing the procedural material and a differentiable version DiffMat in Sec. 3.1. We then provide the detailed descriptions of these three steps in Sec. 3.2, Sec. 3.3 and Sec. 3.4.
3.1 Preliminary
Procedural materials. Procedural material (e.g., Substance materials [1]) is a standard way of modeling spatially-varying materials in the 3D design industry, which generates texture maps algorithmically. Different from the static, pre-drawn or pre-baked texture maps, procedural materials are represented as directed acyclic node graphs, primarily consisting of two types of nodes: generators and filters. Generator nodes create spatial textures from scratch, and filter nodes manipulate input textures using specific image processing operations. Specifically, each node has specific parameters, and changing these parameter values will produce different texture maps, including base color, roughness, normal and other maps. Procedural materials provide many beneficial properties, such as being resolution-independent, re-editability, the ability to be dynamically changed during runtime, more efficient in terms of storage space and memory usage and so on.
DiffMat. Shi et al. developed the DiffMat library [35], skillfully utilizing convolutional neural networks to implement the functionality of most filter nodes, thereby making procedural materials differentiable. Recently, DiffMatV2 [24] was proposed to implement differentiable generator nodes and expand the scope of optimizable parameters of filter nodes. In this work, we utilize DiffMatV2 to optimize the parameters $\theta$ in the node graphs.
@@ 3.2 Multi-view Reference Images Generation
Camera views selection. Since the input mesh is assumed to be segmented beforehand, the mask $M_{R}$ for each material part in a certain camera view can be rendered directly through the renderer. We first adaptively select some camera views to ensure that the number of pixels for each material part exceeds a certain threshold. Note that the front view has to be selected because the reference image of this view generated by Stable Diffusion [34] is more realistic. The number of selected views is unconstrained. For more details about the camera selection strategy, please refer to our supplementary document.
Reference images generation. We use the differentiable renderer PSDR-CUDA [41] to render the normal maps and depth maps of the mesh under the selected camera views. Subsequently, we use Stable Diffusion and Controlnet [43] to generate the reference image of front view under the condition of the corresponding normal map. For other camera views, we also use the reference image of front view as an additional condition to generate reference images. Nonetheless, there may still be slight differences between images of different views. To deal with this problem, we select only one reference image for each material part according to the number of pixels in each view. Moreover, if the image quality is not good, the rendered depth maps will be added as an extra condition.
@@ 3.3 Materials Initialization with Agent Guidance
Figure 3: Material Agent. An example of a chair backrest is shown here. By providing both the image prompt and text prompt to our defined material agent, it can predict all possible material categories in order of likelihood.
@@
However, we find that the classical material classification methods [31, 15] suffer from the baked lighting in generated images and confusion of different materials with similar color distributions. Thus, we design a novel material agent using GPT-4V for accurate material initialization via part-level understanding and object-aware material reasoning, e.g., pillows should be assigned fabric or leather materials instead of metal or wood. Specifically, as shown in Fig. 3, given the masked prompt image and the text prompts, the material agent will provide all possible material types in the order of likelihood. Additionally, Fig. 4 presents some examples of agent response and Fig. 10 illustrates the complete classification process and results for a bed mesh.
@@
Figure 4: Agent Response Examples. On the left are the prompt images provided to the material agent, and on the right are the responses returned by the agent.
@@
Masks segmentation and alignment. Since the object geometry in the images generated by Stable Diffusion is not well aligned with the rendered images, the mask $M_{R}$ on the rendered images can not be directly applied to the reference images. Therefore, we use Matcher [27] to get the mask $M_{I}$ on the reference images. Specifically, as shown in Fig. 2, for camera view $i$, we provide three inputs to Matcher: the rendering of the object textured by white material under specific ambient lighting, the mask $M_{R}$ of a material part on the rendered image, and the reference image. The Matcher then determines the corresponding mask $M_{I}$ of the material part on the reference image. Next, we use median filtering, morphological operations, and region area filtering to smooth $M_{I}$ obtained by Matcher. The final mask $M_{ij}$ of the material part $j$ under camera view $i$ is obtained by $M_{ij}=M_{R}*M_{I}$. $M_{ij}$ will be used on both rendered images and reference images.
@@
Material agent prompt design. For each material part, we blend its mask $M_{ij}$ with the reference image, and then concatenate it horizontally with the reference image to create the prompt image. Given the name of the mesh (e.g., chair), a specific prompt text can be derived from our carefully designed prompt template to guide the material agent. Note that all material parts share the same prompt text, rather than deriving a prompt for every material part. Our template starts by using the GPT-4V system prompt obtained through reverse prompt engineering to enhance performance. Then, we provide a detailed definition of the task along with an example for in-context learning [29]. Subsequently, we employ various prompt techniques to construct the prompt for part-level understanding and object-aware material reasoning. Please refer to our supplementary for the complete prompt template.
@@
Material agent part-level understanding. Similar to the idea of chain of thought [4], we first guide the agent to identify what object is in the marked area, instead of directly inferring the materials. This way allows object information to be taken into account during material reasoning, enabling full use of the extensive world knowledge stored in LLMs. In this part of the prompt, we first explain the prompt image and ask the agent to zoom in on the corresponding area for better recognition. Additionally, we highlight some considerations, such as occlusion issues, and then ask the agent to perform object recognition step-by-step.
@@
Material agent object-aware material reasoning. In this part, we ask the agent to combine pixel information and object information to reason about all possible material types and give prediction results in the order of possibility. We have divided the materials in the DiffMat library into 19 categories including asphalt, bricks, ceramic, concrete, fabric, etc., and require the agent to only choose from these types. We provide the detailed output format and an example for in-context learning [29].
@@
Material matching based on weighted VGG loss. Then we design a multi-scale weighted material matching method based on VGG loss [36], considering the order of possibility. Specifically, we first use the mask $M_{ij}$ of material part $j$ to obtain its bounding box, the cropped mask $M_{ij}^{\prime}$, and the cropped image $C_{ij}$ from the reference image $I_{i}$. Next, we randomly sample a certain number of rectangles $E_{k}^{s}$ on the rendered exemplars at different scales (128×128, 256×256, 512×512) of all materials from the predicted possible types, according to the bounding box size. Then, we apply the cropped mask $M_{ij}^{\prime}$ to these sampled rectangles to obtain masked exemplar rectangles, thereby mitigating the effects of scale or spatial transformations on the texture patterns. Then we assign the initial material $\mathcal{G}_{j}^{0}$ through:
@@
$\mathcal{G}_{j}^{0}=\underset{\mathcal{G}}{arg~{}min}~{}~{}\alpha^{O(\mathcal{G})-1}~{}\frac{1}{K}\sum_{s=1}^{3}\sum_{k=1}^{K}\sum_{l}\sum_{x}\\ \quad(\mathcal{F}_{vgg}^{l}(C_{ij})-\mathcal{F}_{vgg}^{l}(M_{ij}^{\prime}*E_{k}^{s}))^{2},$ (1)
@@
where $\mathcal{G}$ is a procedural material, $\alpha^{O(\mathcal{G})-1}$ is the weight related to the possibility order $O(\mathcal{G})$ of the type of material $\mathcal{G}$ ($O(\mathcal{G})$ starts from 1, $\alpha$ is set to $\frac{5}{3}$), $K$ is the number of sampled rectangles at each scale, $s$ refers to different scales ($s=1$ is 128×128, $s=2$ is 256×256, $s=3$ is 512×512), $\mathcal{F}_{vgg}^{l}(\cdot)$ refers to the normalized VGG feature map extracted from layer $l$, $\sum_{x}$ denotes the sum over pixels $x$, and $E_{k}^{s}$ is the sampled rectangle on the rendered exemplar of material $\mathcal{G}$ at scale $s$. To intuitively understand Eq. 1, we present an example in Fig. 5.
@@
Figure 5: Material Matching Example. The upper part displays the cropped image $C_{ij}$ and the cropped mask $M_{ij}^{\prime}$ obtained through the bounding box. The lower part depicts the masked exemplar rectangles $M_{ij}^{\prime}*E_{k}^{s}$ randomly sampled on the rendered exemplars at different scales of a fabric material.

### v2

@@
[…]
, runauthor = Ziqiang Dang et al., abstract = In this paper, we present TexPro, a novel method for high-fidelity material generation for input 3D meshes given text prompts. Unlike existing text-conditioned texture generation methods that typically generate RGB textures with baked lighting, TexPro is able to produce diverse texture maps via procedural material modeling, which enables physically-based rendering, relighting, and additional benefits inherent to procedural materials. Specifically, we first generate multi-view reference images given the input textual prompt by employing the latest text-to-image model. We then derive texture maps through rendering-based optimization with recent differentiable procedural materials. To this end, we design several techniques to handle the misalignment between the generated multi-view images and 3D meshes, and introduce a novel material agent that enhances material classification and matching by exploring both part-level understanding and object-aware material reasoning. Experiments demonstrate the superiority of the proposed method over existing SOTAs, and its capability of relighting. , keywords = 3D asset creation, text-guided texturing, procedural materials, MLLM, copyright = The Author(s),
[…]
222$\quad$2 Pico, ByteDance Inc, shanghai, 200235, China. E-mail: Bangbang Yang, ybbbbt@gmail.com; Yuewen Ma, mayuewen@bytedance.com.
333$\quad$3 Communication University of Zhejiang, Hangzhou, 310018, China. E-mail: Liang Li, liliang@cuz.edu.cn.
$\quad$ Authors contributed equally.
Manuscript received: 2025-01-01; accepted: 2025-01-01
@@ 1 Introduction
3D mesh texturing is a critical process in the realm of 3D modeling and visualization, which holds significant importance across various applications such as AR/VR, gaming, filmmaking, etc. By transforming a basic geometric shape into a visually detailed and realistic object, mesh texturing not only improves the aesthetic appeal of the model but also greatly enhances immersion and realism in digital environments. To achieve photorealistic rendering that responds accurately to varied lighting conditions, we need to generate object material that includes a set of texture maps for diffuse and specular reflections, glossiness, surface roughness, etc. Thus, traditional mesh texturing typically involves substantial manual effort. Recent advances in 2D image generation have spurred the development of automated mesh texturing methods [1, 2], but they normally generate RGB textures with baked lighting, which limits their applicability in downstream tasks.
@@
In this paper, we propose a novel method for producing high-quality materials for 3D meshes based on user-provided text prompts as shown in Fig. 1. Unlike previous methods that directly generate pre-baked 2D texture maps, our approach is able to produce diverse texture maps via procedural material modeling, which enables physically-based rendering, relighting, and additional benefits inherent to procedural materials, such as resolution independence (i.e., the ability to producing textures at any desired resolution), re-editability, efficient storage, and so on. The proposed system could simplify the workflow for amateur artists, eliminating the need to spend time on tedious parameter tuning for procedural materials. Moreover, the system’s output can serve as useful initialization, allowing artists to further adjust the parameters of nodes in the procedural material graph to achieve the desired detailed effects without starting from scratch.
@@
Specifically, our approach first generates multi-view reference images given the input textual prompt by employing the latest text-to-image models [3, 4], and then obtains diverse texture maps through rendering-based optimization with recent differentiable procedural materials [5, 6]. After the optimization, we can essentially obtain procedural materials for the mesh, thus supporting downstream applications. However, it is non-trivial to design such a system.
@@
First, despite using additional rendered depth or normal maps as conditions [4], the generated images from the text-to-image model (e.g., Stable Diffusion [3]) often lack accurate alignment with the meshes and may not maintain consistency across different views. To handle these misalignment challenges, as shown in Fig. 2, we propose to exploit an improved segmentation method, Matcher [7], to extract the masks of corresponding objects in generated images, which will be further smoothed and integrated with the rendered masks to obtain the final aligned masks. Additionally, we design an adaptive camera sampling method that ensures each material part exceeds a specified pixel count threshold and is assigned a unique material to address the inconsistency across views.
@@
Second, an initial estimation of material properties is requisite for differentiable procedural material rendering. However, the presence of baked lighting in generated images and the confusion of different materials with similar color distributions challenge initial material classification and matching for classical methods [8, 9]. To overcome these challenges, we leverage the capabilities of multi-modal large language models (MLLMs) to create a novel material agent. Given the masked prompt image, the material agent is able to provide the possible material types in the order of likelihood for the corresponding object via part-level understanding and object-aware material reasoning, which further facilitates the rendering-based optimization with procedural materials.
@@
Additionally, to better recover the material properties, we also consider the environmental lighting during optimization. To make the optimization tractable, we adopt a plain lighting setup to fit the lighting environment in the image. Given the object mesh, our method adaptively creates the floor and four walls in the scene of the differentiable renderer, and adaptively places initial area lights on each wall and above the object. Then we optimize the RGB intensity of these lights together with procedural materials.
@@
  • We propose a novel framework that is able to create high-fidelity textures for input meshes given text prompts and enables physically-based rendering (PBR) and relighting.
  • In order to handle the misalignment between the generated multi-view images and 3D meshes, we propose to combine the rendered mask with an advanced segmentation method, Matcher [7], to achieve accurately aligned part-level masks, and design an adaptive camera sampling strategy to deal with inconsistency across views.
@@
  • The experiments show that the proposed method outperforms the existing state of the art, and maintains the capabilities of relighting.
@@ 2 Related Work
2.1 Texture Generation
Texture maps are essential for 3D geometric models. They involve mapping pre-defined 2D planar images onto 3D models, endowing 3D models with vital color information. Traditional texture creation involves cumbersome manual drawing, assembling repetitive patterns, or stitching multi-view images [10, 11]. With the development of increasingly high-quality 3D datasets and advances in text-to-image generative techniques [12, 3, 4], learning-based approaches have been proposed to generate high-quality textures [13, 14, 1, 2, 15]. Text2Tex [1] incorporates a depth-aware image inpainting model to synthesize high-resolution partial textures progressively from multiple viewpoints, which has also been widely employed in text-to-scene tasks [16, 17, 18]. Similarly, TEXTure [2] also applies an iterative scheme to paint 3D models, while presenting a trimap representation and a novel elaborate diffusion sampling process to tackle inconsistency issues. Although existing methods are capable of generating high-quality textures, they are limited to generating only diffuse maps with baked lighting, lacking support for relighting. Additionally, the generated textures often suffer from the Janus (multi-faced) problem [19] and lack realism.
2.2 Procedural Material Modeling and Capture
Procedural material has been the standard method of material modeling in industry. Materials are represented as node graphs denoted with simple image processing operations which are combined to produce real-world spatially varying BRDF [20] material maps. Given user-specified images or text prompts, many approaches [21, 5, 22, 23, 6] have been proposed to recover the parameters of a predefined procedural material graph in agreement with the input. Moreover, Hu et al. [24] moved beyond parameter regression, utilizing the diffusion model to generate graph structures from text or image conditions. It is worth noting that MATch [5] proposes a differentiable procedural material method that translates procedural graphs into differentiable node graphs, enabling single-shot high-quality procedural material capture.
2.3 Differentiable Rendering
Differentiable rendering has consistently remained a significant research topic in computer graphics and vision. The gradient in rendering is required with respect to numerous factors, and several specifically differentiable renderers have been developed and widely used in the community, like Redner [25], Mitsuba 2/3 [26, 27], Nvdiffrast [28] and PSDR-CUDA [29], the aim being to address challenges arising from the non-differentiable terms in the rendering integral. We make use of differentiable rendering to backpropagate the gradient from pixel space to the parameters of procedural materials, thereby optimizing the materials via gradient descent.
@@ 3 Method
3.1 Aims and Approach
Most 3D meshes in online resource repositories like BlenderKit or Sketchfab, as well as meshes from existing 3D datasets [30, 31, 32, 33, 34], have been segmented into different material parts by their authors during creation, e.g., a chair consists of the frame, backrest and cushion. Moreover, part-level segmentation of a complete mesh can be achieved through shape analysis methods [35, 36, 37]. Thus, given textual prompts, we aim to generate appropriate materials for each material part of the input 3D mesh model and achieve photorealistic and relightable texturing.
@@ 3.1 Aims and Approach
As shown in Fig. 2, our method textures the mesh with PBR materials through three steps: multi-view reference images generation, materials initialization under material agent guidance, and materials optimization using differentiable rendering. In the rest of this section, we first introduce procedural materials and a differentiable version DiffMat in Sec. 3.2. We then provide detailed descriptions of these three steps in Sec. 3.3, Sec. 3.4 and Sec. 3.5.
3.2 Preliminaries
Procedural materials. Procedural material (e.g., Substance materials [38]) is a standard way of modeling spatially-varying materials in the 3D design industry, which generates texture maps algorithmically. Different from the static, pre-drawn or pre-baked texture maps, procedural materials are represented as directed acyclic node graphs, primarily consisting of two types of nodes: generators and filters. Generator nodes create spatial textures from scratch, and filter nodes manipulate input textures using certain image processing operations. Specifically, each node has specific parameters, and changing these parameter values will produce different texture maps, including base color, roughness, normal and other maps. Procedural materials provide many beneficial properties, such as resolution-independence, re-editability, the ability to be dynamically changed during runtime, and greater efficiency in terms of storage and memory usage.
DiffMat. Shi et al. developed the DiffMat library [5], skillfully utilizing convolutional neural networks to implement the functionality of most filter nodes, thereby making procedural materials differentiable. Recently, DiffMatV2 [6] was proposed to implement differentiable generator nodes and expand the scope of optimizable parameters of filter nodes. In this work, we utilize DiffMatV2 to optimize the parameters $\theta$ of the node graphs.
@@ 3.3 Multi-view Reference Images Generation
Camera views selection. Since the input mesh is assumed to be segmented beforehand, the mask $M_{R}$ for each material part in a certain camera view can be rendered directly through the renderer. We first adaptively select some camera views to ensure that the number of pixels for each material part exceeds a certain threshold. Note that the front view has to be selected because the reference image of this view generated by Stable Diffusion [3] is more realistic. The number of selected views is unconstrained. For more details of the camera selection strategy, please refer to the appendix.
Reference images generation. We use the differentiable renderer PSDR-CUDA [29] to render the normal maps and depth maps of the mesh under the selected camera views. Subsequently, we use Stable Diffusion and Controlnet [4] to generate the reference image of the front view under the condition of the corresponding normal map. For other camera views, we also use the reference image of the front view as an additional condition to generate reference images. Nonetheless, there may still be slight differences between images of different views. To deal with this problem, we select only one reference image for each material part according to the number of pixels in each view. Moreover, if the image quality is poor, the rendered depth maps are added as an extra condition.
@@ 3.4 Materials Initialization with Agent Guidance
Figure 3: Material Agent. An example of a chair backrest is shown here. Providing both the image prompt and text prompt to our defined material agent allows it to predict all possible material categories in order of likelihood.
@@
However, we find that classical material classification methods [8, 9] suffer from the baked lighting in generated images and confuse different materials with similar color distributions. Thus, we design a novel material agent using GPT-4V for accurate material initialization via part-level understanding and object-aware material reasoning: for example, pillows should be assigned fabric or leather materials, not metal or wood. Specifically, as shown in Fig. 3, given the masked prompt image and the text prompts, the material agent will provide all possible material types in the order of likelihood. Additionally, Fig. 4 presents some examples of agent response and Fig. 10 illustrates the complete classification process and results for a bed mesh.
@@
Figure 4: Agent Response Examples. Left: prompt images provided to the material agent. Right: responses returned by the agent.
@@
Mask segmentation and alignment. Since the object geometry in the images generated by Stable Diffusion is not well aligned with the rendered images, the mask $M_{R}$ on the rendered images cannot be directly applied to the reference images. Therefore, we use Matcher [7] to get the mask $M_{I}$ for the reference images. Specifically, as shown in Fig. 2, for camera view $i$, we provide three inputs to Matcher: the rendering of the object textured using white material under specific ambient lighting, the mask $M_{R}$ of a material part on the rendered image, and the reference image. Matcher then determines the corresponding mask $M_{I}$ of the material part on the reference image. Next, we use median filtering, morphological operations, and region area filtering to smooth $M_{I}$ obtained by Matcher. The final mask $M_{ij}$ of material part $j$ under camera view $i$ is obtained by $M_{ij}=M_{R}*M_{I}$. $M_{ij}$ is used on both rendered images and reference images.
@@
Material agent prompt design. For each material part, we blend its mask $M_{ij}$ with the reference image, and then concatenate it horizontally with the reference image to create the prompt image. Given the name of the mesh (e.g., chair), a specific prompt text is derived from our carefully designed prompt template to guide the material agent. Note that the same prompt text is shared across all material parts, instead of using a separate one for each. Our template starts by using the GPT-4V system prompt obtained through reverse prompt engineering to enhance performance. Then, we provide a detailed definition of the task along with an example for in-context learning [39]. Subsequently, we employ various prompt techniques to construct the prompt for part-level understanding and object-aware material reasoning. Please refer to the appendix for the complete prompt template.
@@
Material agent part-level understanding. Similar to the idea of chain of thought [40], we first guide the agent to identify what object is in the marked area, instead of directly inferring the materials. This allows object information to be taken into account during material reasoning, enabling full use of the extensive world knowledge stored in LLMs. In this part of the prompt, we first explain the prompt image and ask the agent to zoom in on the corresponding area for better recognition. Additionally, we highlight some considerations, such as occlusion issues, and then ask the agent to perform object recognition step-by-step.
@@
Material agent object-aware material reasoning. Next, we ask the agent to combine pixel information and object information to reason about all possible material types and to predict results in the order of probability. We have divided the materials in the DiffMat library into 19 categories including asphalt, bricks, ceramics, concrete, fabric, etc., and require the agent to only choose from these types. We provide the detailed output format and an example for in-context learning [39].
@@
Material matching based on weighted VGG loss. Then we design a multi-scale weighted material matching method based on VGG loss [41], considering the order of likelihood. Specifically, we first use the mask $M_{ij}$ of material part $j$ to obtain its bounding box, the cropped mask $M_{ij}^{\prime}$, and the cropped image $C_{ij}$ from the reference image $I_{i}$. Next, we randomly sample a certain number of rectangles $E_{k}^{s}$ on the rendered exemplars at different scales ($128\times 128$, $256\times 256$, $512\times 512$) for all materials of the predicted possible types, according to the bounding box size. Then, we apply the cropped mask $M_{ij}^{\prime}$ to these sampled rectangles to obtain masked exemplar rectangles, thereby mitigating the effects of scale or spatial transformations on the texture patterns. Then we assign the initial material $\mathcal{G}_{j}^{0}$ through:
@@
$\mathcal{G}_{j}^{0}=\underset{\mathcal{G}}{\text{argmin}}\,\alpha^{O(\mathcal{G})-1}~{}\frac{1}{K}\sum_{s=1}^{3}\sum_{k=1}^{K}\sum_{l}\sum_{x}\\ \quad(\mathcal{F}_{\mathrm{vgg}}^{l}(C_{ij})-\mathcal{F}_{\mathrm{vgg}}^{l}(M_{ij}^{\prime}*E_{k}^{s}))^{2},$ (1)
@@
where $\mathcal{G}$ is a procedural material, $\alpha^{O(\mathcal{G})-1}$ is the weight related to the likelihood order $O(\mathcal{G})$ of the type of material $\mathcal{G}$ ($O(\mathcal{G})$ starts from 1, $\alpha$ is set to $\frac{5}{3}$), $K$ is the number of sampled rectangles at each scale, $s$ refers to different scales ($s=1$ is $128\times 128$, $s=2$ is $256\times 256$, $s=3$ is $512\times 512$), $\mathcal{F}_{\mathrm{vgg}}^{l}(\cdot)$ refers to the normalized VGG feature map extracted from layer $l$, $\sum_{x}$ denotes the sum over pixels $x$, and $E_{k}^{s}$ is the sampled rectangle on the rendered exemplar of material $\mathcal{G}$ at scale $s$. To aid intuitive understanding of Eq. 1, we present an example in Fig. 5.
@@
Figure 5: Material Matching Example. Above: the cropped image $C_{ij}$ and the cropped mask $M_{ij}^{\prime}$ obtained through the bounding box. Below: the masked exemplar rectangles $M_{ij}^{\prime}*E_{k}^{s}$ randomly sampled on the rendered exemplars at different scales of a fabric material.